# Is analogy enough to draw novel adjective-noun inferences?

### v1

@@ 1 Introduction
This paper questions these conclusions, and investigates whether this task can in fact be solved by analogical reasoning, without composition. For example, for counterfeit scarf, one might reason (as in Figure 1): “Is a counterfeit scarf still a scarf? A scarf is an accessory like a watch or a purse, and a counterfeit watch is still a watch, and a counterfeit purse is still a purse, so a counterfeit scarf is most likely still a scarf”. This skips the compositional step of combining the meanings of the words to derive the meaning of the bigram.
We investigate analogical reasoning through two complementary approaches. First, we build a computational model of analogical reasoning which attempts to derive ratings for the low-frequency and zero-frequency (assumed novel) bigrams in the dataset of Ross et al. (2025), by analogy to the high-frequency ones. A computational model allows us to precisely define what we mean by analogy, and explore the consequences of different implementation decisions. Second, we ask human participants to reason analogically, guided by examples and their own intuition of what analogy means. We then evaluate how often they can produce an analogy, and whether the resulting rating distributions derived analogically is the same as the distributions from Ross et al. (2025), where no instructions on how to reason were given. We find that the ratings derived by analogy significantly differ for several bigrams, suggesting that the original participants did not derive (all) their ratings by analogy.
Between the two methods, we find convincing evidence that while analogical reasoning produces similar results in many cases, it is not sufficient to derive the full set of inference data. Thus, we find support for the view that humans must have access to a compositional mechanism. Further, our analogy model performs worse on novel bigrams than the best LLM in Ross et al. (2024), and our analogy model’s successes and failures correlate poorly with those of the best LLM. This suggests that the LLM is not (just) using analogy in the cases where it can generalize, and supports the claim in Ross et al. (2024) that such LLMs are performing some kind of composition (productively combining the meaning of adjective and noun) in these cases.
@@ 2 Related Work
Privative adjectives such as fake pose a challenge for compositional accounts of semantics, since they cannot be simply intersected with the noun (Kamp and Partee, 1995). Multiple accounts have been proposed for how composition with privative adjectives should work (Partee, 2010; del Pinal, 2015; Martin, 2022; Guerrini, 2024 i.a.). Ross et al. (2025) gather a large quantity of offline human judgments on (privative) adjectives and their membership inferences, discussed further in Section 3, and Ross et al. (2024) extend this dataset to assess LLMs. While Ross et al. (2024) do propose a simple analogy baseline to compare to their LLMs, we propose an improved, more powerful and configurable analogy model and present a detailed analysis of its performance.
@@ 4.4 Discussion: Humans
As an alternative metric, we conduct per-bigram Kolmogorov-Smirnoff tests (Holm-Bonferroni adjusted) comparing the distributions predicted by the analogy model to the human distributions. We find that with memorization of the training set, 10 of the predicted distributions are significantly different ($p<0.05$), of which 3 are zero-frequency bigrams; without memorization, this rises to 20. Since we only have a sample size of $n=12$, this is a conservative estimate. Figure 8 in Appendix C.3 shows a selection of such distributions. The fact that the analogy model significantly deviates from the correct distribution for 9–14% of cases supports our conclusion that while analogy is successful in many cases, it does not offer a full explanation.
@@ 5.1 Method
We recruited 176 native American English speakers777 See Appendix B for detailed criteria. on Prolific, of which we excluded 33 for not meeting our native speaker criteria, failed attention checks, or failing to adequately follow our instructions for analogical reasoning (verified based on manual inspection and regular expression searches on the free text entry fields).
@@ 5.2 Results
Figure 5 shows statistics for the types of analogy drawn. We find that 58.4% of analogies use the same adjective as the original bigram, such as knockoff watch $\rightarrow$ knockoff purse, while only 10% change the adjective and use the same noun, such as homemade money $\rightarrow$ counterfeit money. A further 6.2% of analogies use a single noun. While a number of these examples seem intended as same-adjective (such as tiny bed $\rightarrow$ chair), we do see some interesting cases such as artificial rumor $\rightarrow$ lie, which may not be an analogy in the strict sense but are still solving the task by mapping to a known phrase. The remaining 25.4% use a different adjective/modifier and noun.
@@
We also conduct Kolmogorov-Smirnoff tests per-bigram (with Holm-Bonferroni adjustment) to determine which of the distributions are significantly different. Since our $n$ per bigram is quite small for statistical purposes (at best $n=12$, lower if not all participants found an analogy for the bigram), no bigrams are significantly different. We cannot conclude from this that the distributions are indeed the same when analogy is used; the sample size is just too limited. Instead, we plot the distributions for 6 bigrams with the highest JS divergences in Figure 6. The divergence for homemade currency and homemade money (and to a lesser extent tiny proposal) is particularly striking: analogy leads people to dramatically different inferences in these cases, since most homemade and tiny items still clearly qualify as that item.
@@ 6 Conclusion
We find that while many of the novel bigrams in the dataset can indeed be handled successfully by analogy, analogy is not sufficient to explain human behavior fully. Our analogy model diverges from human distributions on at least 9% of the dataset and shows insufficient generalization to zero-frequency bigrams with a JS divergence of 0.25 from humans. Humans both struggle to come up with analogies for 24% of bigrams tested and are led astray when they do for at least 6 bigrams, such as homemade currency. We thus conclude that analogical reasoning is a successful strategy for generalization in a remarkable proportion of the dataset of Ross et al. (2025), but analogy does not suffice to handle the full data. Thus, their conclusion that some mechanism of composition seems necessary to handle the whole range, homemade currency and all, is supported—even if humans need not (and judging by our data, quite possibly do not) invoke it in every case.

### v2

@@ 1 Introduction
This paper questions these conclusions, and investigates whether this task can in fact be solved by analogical reasoning, without composition. For example, for counterfeit scarf, one might reason (as in Figure 1): “Is a counterfeit scarf still a scarf? A scarf is an accessory like a watch or a purse, and a counterfeit watch is still a watch, and a counterfeit purse is still a purse, so a counterfeit scarf is most likely still a scarf”. This skips the compositional step of combining the meanings of the words to derive the meaning of the bigram and further violates the principle of compositionality as stated by Szabó (2012) by referring to information beyond the meaning of the bigram’s parts, namely the inferences associated with other adjective-noun bigrams.
We investigate analogical reasoning through two complementary approaches. First, we build a computational model of analogical reasoning which attempts to derive ratings for the low-frequency and zero-frequency (assumed novel) bigrams in the dataset of Ross et al. (2025), by analogy to the high-frequency ones. A computational model allows us to precisely define what we mean by analogy, and explore the consequences of different implementation decisions. Second, we ask human participants to reason analogically, guided by examples and their own intuition of what analogy means. We then evaluate how often they can produce an analogy, and whether the resulting rating distributions derived analogically are the same as the distributions from Ross et al. (2025), where no instructions on how to reason were given. We find that the ratings derived by analogy significantly differ for several bigrams, suggesting that the original participants did not derive (all) their ratings by analogy.
Between the two methods, we find convincing evidence that while analogical reasoning produces similar results in many cases, it is not sufficient to derive the full set of inference data. Thus, we find support for the view that humans must have access to a compositional mechanism. Further, our analogy model performs worse on novel bigrams than the best LLM in Ross et al. (2024), and our analogy model’s successes and failures correlate poorly with those of the best LLM. This suggests that the LLM is not (just) using analogy in the cases where it can generalize, and supports the claim in Ross et al. (2024) that such LLMs are performing some kind of composition (productively combining the meaning of adjective and noun) in these cases. We share our code and data on GitHub. [^1] https://github.com/rossh2/artificial-intelligence/
@@ 2 Related Work
So-called “privative” adjectives such as fake pose a challenge for compositional accounts of semantics, since they cannot be simply intersected with the noun (Kamp and Partee, 1995). Multiple accounts have been proposed for how composition with privative adjectives should work (Partee, 2010; del Pinal, 2015; Martin, 2022; Guerrini, 2024 i.a.).
Most previous computational work on adjective-noun composition using distributional semantics does not discuss privative adjectives (Baroni and Zamparelli, 2010; Vecchi et al., 2017; Hartung et al., 2017). Boleda et al. (2013) cover 16 “non-intensional” adjectives, including two which are commonly taken to be privative (former, mock; see Nayak et al. (2014) for a classification). Boleda et al. build distributional semantic models of adjective-noun composition that use vector addition and matrix multiplication to model adjective-noun composition, but they do not cover analogy. Cappelle et al. (2018) study the distributional semantics of fake and bigrams in which it occurs, but do not implement any method of composition or generalization.
Ross et al. (2025) gather a large quantity of offline human judgments on (privative) adjectives and their membership inferences, discussed further in Section 3, and Ross et al. (2024) extend this dataset to assess LLMs. While Ross et al. (2024) do propose a simple analogy baseline to compare to their LLMs, we propose an improved, more powerful and configurable analogy model and present a detailed analysis of its performance.
@@ 4.4 Discussion: Humans
As an alternative metric, we conduct per-bigram Kolmogorov-Smirnoff tests (Holm-Bonferroni adjusted) comparing the distributions predicted by the analogy model to the human distributions. We find that with memorization of the training set, 10 of the predicted distributions are significantly different ($p<0.05$), of which 3 are zero-frequency bigrams; without memorization, this rises to 20. Since we only have a sample size of $n=12$, this is a conservative estimate. Figure 8 in Appendix C.3 shows a selection of such distributions. The fact that the analogy model significantly deviates from the correct distribution for these cases supports our conclusion that while analogy is successful in most cases, it does not offer a full explanation.
@@ 5.1 Method
We recruited 176 native American English speakers888 See Appendix B for detailed criteria. on Prolific, of which we excluded 33 for not meeting our native speaker criteria, failed attention checks, or failing to adequately follow our instructions for analogical reasoning (verified based on manual inspection and regular expression searches on the free text entry fields).
@@ 5.2 Results
Figure 5 shows statistics for the types of analogy drawn. We find that 58.4% of analogies use the same adjective as the original bigram, such as knockoff watch $\rightarrow$ knockoff purse, while only 10% change the adjective and use the same noun, such as homemade money $\rightarrow$ counterfeit money. A further 6.2% of analogies use a single noun. While a number of these single-noun analogies seem intended as same-adjective analogies (such as tiny bed $\rightarrow$ (tiny) chair), we do see some interesting cases such as artificial rumor $\rightarrow$ lie, which may not be an analogy in the strict sense but are still solving the task by mapping to a known phrase. The remaining 25.4% use a different adjective/modifier and noun.
@@
We also conduct Kolmogorov-Smirnoff tests per-bigram (with Holm-Bonferroni adjustment) to determine which of the distributions are significantly different. Since our $n$ per bigram is quite small for statistical purposes (at best $n=12$, lower if not all participants found an analogy for the bigram), no bigrams are significantly different. We cannot conclude from this that the distributions are indeed the same when analogy is used; the sample size is just too limited. Instead, we plot the distributions for 6 bigrams with the highest JS divergences in Figure 6. The divergence for homemade currency and homemade money (and to a lesser extent false friend) is particularly striking: analogy leads people to dramatically different inferences in these cases, since most homemade and many false items (such as false rumor) still clearly qualify as an instance of the noun.
@@ 6 Conclusion
We find that while many of the novel bigrams in the dataset can indeed be handled successfully by analogy, analogy is not sufficient to explain human behavior fully. Our analogy model diverges significantly from human distributions on 20 bigrams and shows insufficient generalization to zero-frequency bigrams, with a JS divergence of 0.25 from humans. Humans both struggle to come up with analogies for 24% of bigrams tested and are led astray when they do for several bigrams, such as homemade currency. We thus conclude that analogical reasoning is a successful strategy for generalization in a remarkable proportion of the dataset of Ross et al. (2025), but analogy does not suffice to handle the full data. Thus, their conclusion that some mechanism of composition seems necessary to handle the whole range, homemade currency and all, is supported—even if humans need not (and judging by our data, quite possibly do not) invoke it in every case. This conclusion is similar to the result of Albright and Hayes (2003), who found that an analogical model of English past tense morphology did not explain participant behavior well, and concluded that speakers used abstract rules to generalize rather than analogy.
@@
Standards of evidence for composition
This paper contributes to a broader discussion about the standards of evidence required for composition (McCurdy et al., 2024; Pavlick, 2025). If behavioral experiments about generalization can provide evidence about composition (and not all researchers believe they can), we must be sure to rule out other methods of generalization such as analogy. We further need to ensure we have a precise enough definition of compositionality to capture our intuition that analogy, by virtue of referring to information not (obviously) included in the meanings of the parts, is not a kind of composition (Szabó, 2012). By making an explicit model of analogical reasoning, we can both show the way in which it requires this additional information and show that analogical reasoning fails to generalize in the expected way, relative to our human data.